Testing the ESM TTM table: Checking freq and emgains:

Argos source in the beam: global offsets 3.68 3.33 -1.07 -105 75

Executing tests with 1" seeing files today since typically the most unstable. Upping the frequncy seemed to produce less safe skips with the 1" seeing files

Center star failed to complete although appeared successful in flao gui:

WfsArbitrator_34698|INF|       547|2020-12-30 06:06:57.489523|             MAIN > source_acq.acquire(): starting source acquisition
WfsArbitrator_34698|INF|       548|2020-12-30 06:06:57.491038|             MAIN > Exception: thGetVar wfsarb.L.CLICK_ACQUIRE error(-3006): Variable not found error - Assuming automatic acquisition
WfsArbitrator_34698|INF|       549|2020-12-30 06:07:01.949692|             MAIN > Star found at x=124.0, y=125.8, flux=19149.3
WfsArbitrator_34698|INF|       550|2020-12-30 06:07:01.951382|             MAIN > source_acq.acquire(): star found at X=496.02,Y=503.20
WfsArbitrator_34698|INF|       551|2020-12-30 06:07:01.951444|             MAIN > source_acq.acquire(): current error X=0.063 mm, Y=0.022 mm
WfsArbitrator_34698|INF|       552|2020-12-30 06:07:01.951480|             MAIN > source_acq.acquire(): target reached.

retry was sucessful

Error when closing the loop on 12th mag:
AdSecArbitrator_107988|INF|      8748|2020-12-30 06:42:22.991228|  COMMANDHANDLER > Command not executed (status RETRY) (errstr IDL The disturbance triggered by SC is still running)
AdSecArbitrator_107988|WAR|      8749|2020-12-30 06:42:22.991258|            MAIN > LoadDisturbTag: WARNING - IDL The disturbance triggered by SC is still running
idl_107910         |INF|      1327|2020-12-30 06:42:22.990853|             MAIN > IDL The disturbance triggered by SC is still running
20203012 06:42:34.5661 Alert: HighOrderOffload: HighOrderOffload vector ready

Worked after Closed with disturbance sdisabled but was sticky until I figured that out.

Skip frame detected. WOuld skip afgter running optimize gain even with zero gains applied.

Ref Mag aorgos shutter Seeing Bin Freq emgains tt ho1 ho2 average pupil intensity photons/subapp sloperms TN
9.9 fully s2.0 ovfreq500 1 497 200 6.26 9.83 0.08 100 62 0.757 055515
11.0 68 s2.0 ovfreq500 2 400 600 7.29 12.26 0.08 ~680 136.3 0.665 forgot to record#
11.0 68 s2.0 ovfreq500 2 400 100 7.29 12.26 0.08 ~100 137.0 0.665 062757
12.0 62 s2.0 ovfreq500 2 400 600 2.74 9.91 0.08 ~270 54.8 0.669 065751
12.0 62 s2.0 ovfreq500 2 400 250 2.74 9.91 0.08 ~110 54.8 0.665 070203
13.0 59 s2.0 ovfreq500 2 400 600 9.00 8.57 0.08 ~110 22.7 0.684 071432
13.0 59 s2.0 ovfreq200 2 200 300 1.89 14.78 0.08 ~130 44.7 0.675 072507
13.0 59 s2.0 ovfreq250 4 400 100 0.65 0.55 0.10 ~90 100.2 0.511 75356 Safeskipping 74% during optimization although plots very nice, manually set gains
14.0 55 s2.0 ovfreq250 4 400 600 0.55 0.50 0.10 ~200 37.9 0.547 082000
14.0 55 s2.0 ovfreq250 2 400 600 4.08 10.77 0.08 ~50 9,6 0.760 083322
14.0 55 s2.0 ovfreq250 2 200 600 4.08 10.77 0.08 ~90 16.5 0.760 084409 Garbage, some shape not removed
14.0 55 s2.0 ovfreq200 2 200 600 9.00 1.64 0.08 ~90 17.5 0.673 085710 no disturbance file, the dustrurbance file is putting too much on the mirror
14.0 55 s1.5 ovfreq200 2 400 600 9.00 10.48 0.08 ~50 9.3 0.673 085710 weird shape. We had shifted so possible viegnette, clearing optics and starting again."
14.0 55 s1.5 ovfreq200 2 400 600 4.10 9.24 0.09 ~50 9.3 0.673 092444

Shifting gears to open loop data collection: (closed with zero gains)

Ref Mag Seeing Bin Freq emgains LUCI (luci1.20201230.) FWHM (arcsec) TN
14.0 s2.0 ovfreq500 2 400 600 51-53 0.75 95325
14.0 s2.0 ovfreq200 2 200 600 54-56 0.75 95711
14.0 s2.0 ovfreq200 2 100 600 60-62 0.75 100335 (safeskips about 10%)
14.0 s1.5 ovfreq200 2 100 600 63-65 0.64 100641
14.0 s1.5 ovfreq200 2 200 600 66-68 0.64 100910
14.0 s1.5 ovfreq200 2 400 600 69-71 0.65 101223
14.0 zero dist 2 400 600 72-74 0.5 but psf has an odd shape 101655
14.0 zero dist 2 200 600 75-77 0.27 TN applied 10 modes then bumped back down to zero to remove a bit of the built up shape
14.0 zero dist 2 400 600 78-80 0.30 102339
14.0 zero dist 2 100 600 81-83 0.30 102618
9.9 zero dist 1 500 200 84-86 0.30 103432
9.9 zero dist 1 200 20 87-89 0.30 103657
9.9 zero dist 1 100 20 90-92 0.30 103657
9.9 s1.5 ovfreq200 1 100 20 93-95 0.46, 0.59, 0.50 104203 IQ jumpy? 200Hz already more stable...
9.9 s1.5 ovfreq200 1 200 20 96-98 0.5 104203
9.9 s1.5 ovfreq500 1 500 200 99-101 0.58, 0.58, 0.61 105444
9.9 s1.5 ovfreq500 1 500 200 102-104 0.5, 0.54, 0.54   taking another set
9.9 s2.0 ovfreq500 1 500 200 105-107 0.67, 0.69, 0.72 110324
9.9 s2.0 ovfreq200 1 200 20 108-110 0.75 110324 Alex increased NDITS of his exposures to match teh bin2 times
9.9 s2.0 ovfreq200 1 100 20 111-113 0.67, 0.69, 0.67 112142
10:51 OCAM not live, changing binnings brought it back to life OCAM powered off and ARGOS source parked.

-- JenniferPower - 30 Dec 2020
Topic revision: r3 - 30 Dec 2020, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback